Ruby is three years older than Sam. Sum of their age is 33. How old are both of them?

Found 2 solutions by algebrahouse.com, Lstark01:
Answer by algebrahouse.com(1659)

"Ruby is three years older than Sam. Sum of their age is 33. How old are both of them?"

x = Sam's age now
x + 3 = Ruby's age now

x + x + 3 = 33 {sum of their ages is 33}
2x + 3 = 33 {combined like terms}
2x = 30 {subtracted 3 from both sides}
x = 15 {divided both sides by 2}
x + 3 = 18 {substituted 15, in for x, into x + 3}

Sam is 15 and Ruby is 18
www.algebrahouse.com

Answer by Lstark01(4)

Let R=Ruby, S=Sam. The two equations you will use are R=S+3 and R+S=33. Since R is already solved, use the substitution method and place it in the second equation:
(S+3)+S=33
2S+3=33
-3 on both sides
2S=30
Divide by 2 on both sides
S=15
Then, if you place S=15 into either equation, you can solve for R.
R=15+3
R=18
or
R+15=33
-15 on both sides
R=18